OpenStudy (anonymous):

What is the gravitational potential energy of a book with a mass of 0.30 kg on a 1.0 m high desk?

OpenStudy (anonymous):

use this equation, E=mgh

OpenStudy (anonymous):

Is it 0.3 joules?

OpenStudy (anonymous):

you should multiply it by the g.

OpenStudy (anonymous):

PE(grav) = mass x gravitational acceleration x h PE = 0.30KG x 9.8m.s-1 x 1.0 PE = 2.94J or 2.9J to 2 significant digits

OpenStudy (anonymous):

so its 3 joules

OpenStudy (anonymous):

If the index of refraction of glass is 1.5, calculate the velocity of light in glass. _________ m/s light always travels as 3.00 x 108 m/s 4.5 x 108 5.0x 10-9 2.0 x 108

OpenStudy (anonymous):

approximately, yes

OpenStudy (lastdaywork):

use formula - refractive index = velocity of light in vacuum/velocity of light in medium

OpenStudy (anonymous):

n = c/v c= 3.00 x 10^8 n= 1.5 therefore v= c/n

OpenStudy (anonymous):

2.0 x 10^8 m.s^-1

OpenStudy (anonymous):

oh ok